# A note on thermal effects in non-linear models for plasma-based acceleration

### v1

@@ 4 Outlook and Conclusions
Figure 8: Longitudinal electric fields from PIC simulations (blue) compared with model predictions (dashed orange) for a cold background plasma (top panel) and for a warm plasma with $k_{b}T_{i}=1\;\rm{keV}$ (bottom panel). The model accurately reproduces the accelerating field, with noticeable deviations only at the tail of the bubble, consistent with previous observations [8, 9]. As also reported in [22], the presence of a warm background plasma does not significantly alter the field structure inside the bubble.
We have studied the impact of finite electron temperature on the structure of the first plasma bubble in the blowout regime of plasma wakefield acceleration. Building upon the cold-plasma model introduced by Lu et al. in [8, 9], we developed an extended theoretical framework that incorporates stochastic initial conditions and stochastic ODE coefficients arising from thermal momentum distributions. This extension represents a paradigm shift from the deterministic trajectory $r_{b}(\xi)$ of the cold-plasma case: the resulting ODE of the theory must be solved for an ensemble of particles, and characterized through stochastic averages and fluctuations. This must go together with a re-parametrization of the source term adopted in the model: indeed, as shown by PIC simulations, the region where the source term exhibits a sharp rise from the bubble region to the electron sheath is regularized by a non-zero temperature and requires, for each temperature, novel optimal values of the sheath thickness $\Delta$ that one can identify through an a posteriori parametric matching procedure. With this correction, the model manages to qualitatively reproduce the temperature induced contraction of both the longitudinal and transverse bubble sizes. This work is a first step toward the extension of theoretical models to thermal regimes, and can be improved along several directions. First, the seminal model by Lu et al. in [8, 9] has since been followed by more recent approaches, such as the multi-shell model of Dalichaouch et al. [15] and the energy conservation formulation by Golovanov et al. [17]. One wonders how thermal effects could be implemented in these frameworks, and whether their warm extensions would improve on the current work. Second, the sheath-width parametrization employed here was based on a manual fitting procedure; more refined strategies, including data-driven or machine learning approaches, could be explored to systematically optimize the model and reduce reliance on manual tuning. Third, temperature effects can be modeled at the hydrodynamic scale by choosing an appropriate closure for the fluid equations [33, 34]. One could then explore whether the Lu model can be reconstructed starting from such fluid equations, with thermal effects included from the beginning.

### v2

@@ 4 Outlook and Conclusions
Figure 8: Longitudinal electric fields from PIC simulations (orange) compared with model predictions (dashed blue) for a cold background plasma (top panel) and for a warm plasma with ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k_{B}}T_{i}=1\;\rm{keV}$ (bottom panel).
We have studied the impact of finite electron temperature on the structure of the first plasma bubble in the blowout regime of plasma wakefield acceleration. Building upon the cold-plasma model introduced by Lu et al. in [32, 30], we developed an extended theoretical framework that incorporates stochastic initial conditions and stochastic ODE coefficients arising from thermal momentum distributions. This extension represents a paradigm shift from the deterministic trajectory $r_{b}(\xi)$ of the cold-plasma case: the resulting ODE of the theory must be solved for an ensemble of particles, and characterized through stochastic averages and fluctuations. This must go together with a re-parametrization of the source term adopted in the model: indeed, as shown by PIC simulations, the region where the source term exhibits a sharp rise from the bubble region to the electron sheath is regularized by a non-zero temperature and requires, for each temperature, novel optimal values of the sheath thickness $\Delta$ that one can identify through an a posteriori parametric matching procedure. With this correction, the model manages to qualitatively reproduce the temperature induced contraction of both the longitudinal and transverse bubble sizes. This work is a first step toward the extension of theoretical models to thermal regimes, and can be improved along several directions. First, the seminal model by Lu et al. in [32, 30] has since been followed by more recent approaches, such as the multi-shell model of Dalichaouch et al. [8] and the energy conservation formulation by Golovanov et al. [15]. One wonders how thermal effects could be implemented in these frameworks, and whether their warm extensions would improve on the current work. Second, the sheath-width parametrization employed here was based on a manual fitting procedure; more refined strategies, including data-driven or machine learning approaches, could be explored to systematically optimize the model and reduce reliance on manual tuning. Third, in the presence of temperature effects, one could explore whether part of the parametrization of the electron sheath (e.g., its smoothing and the decrease in the peak) can be predicted from "first principles" starting from the cold data or derived from more continuum models [46, 47].